Human blood circulation system has about 1×10^9 capillary vessels. Each vessel has a diameter 8μm (1μm=1.0×10^(-6) m). Assuming the cardiac output is 4.8 L/min, determine the speed of blood flow through the capillary vessel.
Answer #1

1 L = 10-3 m3

1 min = 60 sec

Flow rate of blood = Q = 4.8 L/min

Converting to m3/s,

107 Q = 4.8 x 60

Q = 8 x 10-5 m3/s

Diameter of the capillary = D = 8 \mum = 8 x 10-6 m

Cross-sectional area of one capillary = A

A = \piD2/4

A = \pi(8x10-6)2/4

A = 5.0265 x 10-11 m2

Number of capillary vessels = n = 1 x 109

Speed of the blood flow through the capillary vessel = V

Q = nAV

8x10-5 = (1x109)(5.0265x10-11)V

V = 1.59 x 10-3 m/s

Speed of blood flow through the capillary vessel = 1.59 x 10-3 m/s